### v1

# Self-avoiding walks on cubic graphs and local transformations

Abstract.
We study self-avoiding walks (SAWs) on infinite quasi-transitive cubic graphs under local transformations that replace each degree-$3$ vertex by a finite, symmetric three-port gadget. To each gadget we associate a two-port SAW generating function $g(x)$, defined by counting SAWs that enter and exit the gadget through prescribed ports. Our first main result shows that, if $G$ is cubic and $G_{1}=\phi(G)$ is obtained by applying the local transformation at every vertex, then the connective constants $\mu(G)$ and $\mu(G_{1})$ satisfy the functional relation
$\mu(G)^{-1}=g\bigl(\mu(G_{1})^{-1}\bigr).$
We next consider critical exponents defined via susceptibility-type series that do not rely on an ambient Euclidean dimension, and prove that the exponents $\gamma$ and $\eta$ are invariant under local transformations; moreover $\nu$ is invariant under a standard regularity hypothesis on SAW counts (a common slowly varying function).
Our second set of results concerns bipartite graphs, where the local transformation is applied to one colour class (or to both classes, possibly with different gadgets). In this setting we obtain an analogous relation
$\mu(G)^{-2}=h\bigl(\mu(G_{\mathrm{e}})^{-1}\bigr),$
with $h(x)=xg(x)$ when only one class is transformed and $h(x)=g_{\phi_{1}}(x)\,g_{\phi_{2}}(x)$ when both are transformed. We further present explicit families of examples, including replacing each degree-3 vertex by a complete-graph gadget $K_{N}$.
Key words and phrases:
Self-avoiding walk, connective constant, cubic graph, Fisher transformation, quasi-transitive graph
2010 Mathematics Subject Classification:
05C30, 82B20
1. Introduction
A self-avoiding walk (SAW) on a graph $G$ is a path that visits no vertex more than once. SAWs were introduced in the chemical theory of polymerization (see Flory [5]) and have since been central objects in probability, combinatorics, and statistical physics; see, for example, Madras and Slade [13]. For an infinite quasi-transitive graph $G$, the number $\sigma_{n}$ of $n$-step SAWs grows exponentially, and the associated exponential growth rate
$\mu=\mu(G)=\lim_{n\to\infty}\sigma_{n}^{1/n}$
is the connective constant of $G$. Exact values of connective constants are known only in a small number of nontrivial cases.
This paper studies how SAWs and their asymptotic parameters behave under local transformations of cubic graphs. Informally, a local transformation replaces every degree-$3$ vertex by a fixed finite gadget with three distinguished outer vertices (ports) attached to the three incident edges, subject to a symmetry condition ensuring that the three ports are indistinguishable from the point of view of the gadget (Definition 3.1). A canonical example is the Fisher transformation, which replaces each vertex by a triangle; Fisher-type transformations play an important role in exact mappings between two-dimensional lattice models (e.g. Ising/dimer/vertex models) and have also been used in the SAW literature.
Connective constants under local transformations. Given a local transformation $\phi$, we associate to the gadget a two-port generating function $g(x)$ (Definition 3.2), defined by counting SAWs that enter the gadget through one port-edge and exit through another, while otherwise staying inside the gadget. Our first main theorem (Theorem 3.3) shows that this local data determines the connective constant of the transformed graph: if $G$ is infinite, connected, quasi-transitive and cubic, and $G_{1}=\phi(G)$ is obtained by applying $\phi$ at every vertex, then
(1.1) $\mu(G)^{-1}=g\bigl(\mu(G_{1})^{-1}\bigr).$
Equivalently, $\mu(G_{1})^{-1}$ is the unique positive solution of $g(x)=\mu(G)^{-1}$. This functional relation may be viewed as a substitution identity for SAW partition functions, and it extends the familiar relation arising from the Fisher transformation.
Critical exponents. Beyond exponential growth, SAWs are expected to exhibit universal power-law corrections, captured by critical exponents. Rather than assuming an ambient Euclidean dimension, we work with graph-theoretic definitions of the exponents $\gamma,\eta,\nu$ via the finiteness of natural SAW generating series (Section 3). Our second main theorem (Theorem 3.4) proves that local transformations preserve these exponents under mild hypotheses: $\gamma$ and $\eta$ are invariant, and $\nu$ is invariant assuming the SAW counts satisfy a standard slowly varying correction (3.7) with a common slowly varying function.
Transforming one colour class of a bipartite graph. Our third main result concerns bipartite graphs. Let $G$ be an infinite, connected, quasi-transitive bipartite graph with colour classes (black/white), and suppose at least one class is cubic. Let $G_{\mathrm{e}}$ be obtained by applying a local transformation to all vertices in one class (or, in the fully cubic case, applying possibly different local transformations to the two classes). Theorem 3.5 shows that the connective constants satisfy
(1.2) $\mu(G)^{-2}=h\bigl(\mu(G_{\mathrm{e}})^{-1}\bigr),$
with $h(x)=xg(x)$ if only one class is transformed, and $h(x)=g_{\phi_{1}}(x)\,g_{\phi_{2}}(x)$ if both classes are transformed. This yields new exact connective constants when combined with known exact values.
The special case in which each cubic vertex is replaced by a triangle is the Fisher transformation. The effect of the Fisher transformation on self-avoiding walks was analyzed by Grimmett and Li [6], who derived an explicit functional relation between the connective constants of $G$ and its Fisher transform and proved invariance of the principal critical exponents under iteration.
The present paper extends [6] from Fisher triangles to general local transformations in which each degree-$3$ vertex is replaced by an arbitrary finite three-port gadget satisfying a natural symmetry condition (Definition 3.1). Under this general framework we obtain corresponding functional relations for connective constants (Theorems 3.3 and 3.5) and prove invariance of the critical exponents under verifiable hypotheses (Theorem 3.4), including the case where the transformation is applied to only one colour class of a bipartite graph (Theorem 3.5). .
Further examples. We include explicit families of gadgets beyond Fisher triangles. In particular, replacing each cubic vertex by a complete graph $K_{N}$ with three designated outer vertices yields an explicit polynomial $g_{N}(x)$ (Example 6.1), and hence an explicit implicit relation between the connective constants of $G$ and $\phi^{(N)}(G)$ for any $N\geq 3$. When $N=3$, this corresponds to Fisher transformation. See Figure 1.1 for local replacements of a degree-3 vertex by a complete graph $K_{3}$ and $K_{4}$.
[…]
Figure $1.$ Local replacement at a degree-$3$ vertex: (a) the original star, (b) Fisher transformation (triangle gadget), and (c) replacement by a $K_{4}$ gadget. See also Definition 3.1 and Figure 3.1 in the draft.
A useful class of examples comes from vertex-transitive TLF-planar cubic graphs. Following Renault ([14]), planarity plus vertex-transitivity imposes strong geometric symmetry on embeddings: the action of graph automorphisms can (locally) be represented by isometries of the ambient geometry (spherical, Euclidean, or hyperbolic).
Let $G$ be a transitive, TLF-planar cubic graph, and let $\langle a,b,c\rangle$ be its type-vector, i.e., the sizes of the three faces incident to a vertex, with the convention $1/\infty=0$. Define
$f(G):=\frac{1}{a}+\frac{1}{b}+\frac{1}{c}.$
By [14] (see also [8]), $G$ admits a planar TLF embedding by regular polygons in the sphere / Euclidean plane / hyperbolic plane according as $f(G)>\frac{1}{2}$, $f(G)=\frac{1}{2}$, or $f(G)<\frac{1}{2}$,
In the Euclidean case $f(G)=\frac{1}{2}$, the only possible type-vectors are $\langle 6,6,6\rangle$, $\langle 3,12,12\rangle$, $\langle 4,8,8\rangle$, $\langle 4,6,12\rangle$, and $\langle 4,4,\infty\rangle$; see Figure 1.2.
$\langle 6,6,6\rangle$ $\langle 4,8,8\rangle$ $\langle 3,12,12\rangle$
$\langle 4,6,12\rangle$ $\langle 4,4,\infty\rangle$
Figure 1.2. Local regular-polygon patches for the five Euclidean type-vectors of transitive TLF-planar cubic graphs.
Organization. Section 2 collects basic notation. Local transformations and the associated generating function $g(x)$ are introduced in Section 3, where we also state Theorems 3.3–3.5. The proofs of Theorems 3.3–3.4 and Theorem 3.5 appear in Sections 4 and 5, respectively. We conclude with explicit examples in Section 6.
2. Notation
All graphs studied henceforth in this paper will be assumed infinite, connected, and simple (in that they have neither loops nor multiple edges). An edge $e$ with endpoints $u$, $v$ is written $e=\langle u,v\rangle$. If $\langle u,v\rangle\in E$, we call $u$ and $v$ adjacent and write $u\sim v$. The degree of vertex $v$ is the number of edges incident to $v$, denoted $\deg(v)$. A graph is called cubic if all vertices have degree $3$. The graph-distance between two vertices $u$, $v$ is the number of edges in the shortest path from $u$ to $v$, denoted $d_{G}(u,v)$.
The automorphism group of the graph $G=(V,E)$ is denoted ${\mathcal{A}}={\mathcal{A}}(G)$. The graph $G$ is called quasi-transitive if there exists a finite subset $W\subseteq V$ such that, for $v\in V$ there exists $\alpha\in{\mathcal{A}}$ such that $\alpha v\in W$. We call such $W$ a fundamental domain, and shall normally (but not invariably) take $W$ to be minimal with this property. The graph is called vertex-transitive (or transitive) if the singleton set $\{v\}$ is a fundamental domain for some (and hence all) $v\in V$.
A walk $w$ on $G$ is an alternating sequence $v_{0}e_{0}v_{1}e_{1}\cdots e_{n-1}v_{n}$ of vertices $v_{i}$ and edges $e_{i}$ such that $e_{i}=\langle v_{i},v_{i+1}\rangle$. We write $|w|=n$ for the length of $w$, that is, the number of edges in $w$.
Let $n\in{\mathbb{N}}$, the natural numbers. A $n$-step self-avoiding walk (SAW) on $G$ is a walk containing $n$ edges that includes no vertex more than once. Let $\sigma_{n}(v)$ be the number of $n$-step SAWs starting at $v\in V$. It was shown by Hammersley [9] that, if $G$ is quasi-transitive, there exists a constant $\mu=\mu(G)$, called the connective constant of $G$, such that
(2.1) $\mu=\lim_{n\to\infty}\sigma_{n}(v)^{1/n},\qquad v\in V.$
It will be convenient to consider also SAWs starting at ‘mid-edges’. We identify the edge $e$ with a point (also denoted $e$) placed at the middle of $e$, and then consider walks that start and end at these mid-edges. Such a walk is self-avoiding if it visits no vertex or mid-edge more than once, and its length is the number of vertices visited.
The minimum of two reals $x$, $y$ is denoted $x\wedge y$, and the maximum $x\vee y$.
3. Local transformation
Definition 3.1 (Local transformation).
Let $G=(V,E)$ be a simple graph, and let $v\in V$ be a vertex of degree $3$. A local transformation $\phi$ acts at $v$ by replacing $v$ with a finite graph (a gadget) $\phi(v)$ that satisfies the following conditions:
  1. (1)
    The gadget $\phi(v)$ comes with three distinguished outer vertices, each of which is attached to exactly one of the three edges incident to $v$ (one outer vertex per incident edge).
  2. (2)
    For any degree-$3$ vertices $v_{1},v_{2}\in V$, the gadgets $\phi(v_{1})$ and $\phi(v_{2})$ are isomorphic (via an isomorphism that sends outer vertices to outer vertices).
  3. (3)
    Let $w_{1},w_{2},w_{3}$ be the three outer vertices of $\phi(v)$. There exists a subgroup of the automorphism group of the gadget that
    • acting transitively on $\{w_{1},w_{2},w_{3}\}$; equivalently, for any $w_{i},w_{j}\in\{w_{1},w_{2},w_{3}\}$ there exists an automorphism of $\phi(v)$ mapping $w_{i}$ to $w_{j}$; and
    • preserving the set of distinguished outer vertices $\{w_{1},w_{2},w_{3}\}$
The resulting graph is denoted by $\phi(G)$.
One special case of local transformations defined above is the so-called Fisher transformation; in which each vertex is replaced by a triangle. This transformation has been valuable in the study of the relations between Ising, dimer, and general vertex models (see [2, 4, 11, 12]), and more recently of SAWs on the Archimedean lattice denoted $(3,12^{2})$ (see [7, 10]).
In the remainder of this paper, we make use of the more general local transformation in the context of SAWs and the connective constant. It will be applied to cubic graphs, of which the hexagonal and square/octagon lattices are examples.
It is convenient to work with graphs with well-defined connective constants, and to this end we assume that $G=(V,E)$ is quasi-transitive and connected, so that its connective constant is given by (2.1). We write $\phi(G)$ for the graph obtained from the cubic graph $G$ by applying the local transformation at every vertex. Obviously the automorphism group of $G$ induces an automorphism subgroup of $\phi(G)$.
Definition 3.2.
Let $G=(V,E)$ be a simple graph. For a degree-$3$ vertex $v\in V$, let $\phi$ be a local transformation as in Definition 3.1. Let $e_{1},e_{2},e_{3}$ be the three edges of $G$ incident to $v$. Define
(3.1) $g(x)\;=g_{\phi}(x)=\;\sum_{\pi\in\Sigma_{\phi(v),e_{1},e_{2}}}x^{|\pi|},$
where $\Sigma_{\phi(v),e_{1},e_{2}}$ denotes the set of all self-avoiding walks (SAWs) that start at the midpoint of $e_{1}$, end at the midpoint of $e_{2}$, and otherwise visit only vertices of the gadget $\phi(v)$. $|\pi|$ denote the total number of vertices in the SAW $\pi$.
By Definition 3.1(2), the gadgets $\phi(v)$ are mutually isomorphic for all degree-33 vertices $v$, so the power series $g(x)$ is independent of the choice of $v$. Moreover, by Definition 3.1(3), the value of $g(x)$ is the same if $\{e_{1},e_{2}\}$ is replaced by any other pair among $\{e_{1},e_{3}\}$ and $\{e_{2},e_{3}\}$ in (3.1).
Theorem 3.3.
Let $G$ be an infinite, quasi-transitive, connected, cubic graph, $G_{1}=\phi(G)$. Then The connective constants $\mu_{1}$ of $G_{1}$ and $\mu$ of $G$ satisfy $\mu^{-1}=g(\mu_{1}^{-1})$ where $g(x)$ is defined as in (3.1).
We turn to the topic of critical exponents, beginning with a general introduction for the case when there exists a periodic, locally finite embedding of $G$ into ${\mathbb{R}}^{d}$ with $d\geq 2$. The case of general $G$ has not not been studied extensively, and most attention has been paid to the hypercubic lattice ${\mathbb{Z}}^{d}$. It is believed (when $d\neq 4$) that there is a power-order correction, in the sense that there exists $A_{v}>0$ and an exponent $\gamma\in{\mathbb{R}}$ such that
(3.2) $\sigma_{n}(v)\sim A_{v}n^{\gamma-1}\mu^{n}\qquad\text{as }n\to\infty,\qquad v\in V.$
Furthermore, the value of the exponent $\gamma$ is believed to depend on $d$ and not further on the choice of graph $G$. When $d=4$, (3.2) should hold with $\gamma=1$ and subject to the inclusion on the right side of the logarithmic correction term $(\log n)^{1/4}$. See [1, 13] for accounts of critical exponents for SAWs.
Let $v\in V$ and
(3.3) $Z_{v,w}(x)=\sum_{n=0}^{\infty}\sigma_{n}(v,w)x^{k},\qquad w\in V,\ x>0,$
where $\sigma_{n}(v,w)$ is the number of $n$-step SAWs with endpoints $v$, $w$. It is known under certain circumstances that the generating functions $Z_{v,w}$ have radius of convergence $\mu^{-1}$ (see [13, Cor. 3.2.6]), and it is believed that there exists an exponent $\eta$ and constants $A^{\prime}_{v}>0$ such that
(3.4) $Z_{v,w}(\mu^{-1})\sim A_{v}^{\prime}{d_{G}(v,w)}^{-(d-2+\eta)}\qquad\text{as }d_{G}(v,w)\to\infty.$
Let $\Sigma_{n}(v)$ be the set of $n$-step SAWs from $v$, and write $\langle\cdot\rangle_{n}^{v}$ for expectation with respect to uniform measure on $\Sigma_{n}(v)$. Let $\|\pi\|$ be the graph-distance between the endpoints of a SAW $\pi$. It is believed (when $d\neq 4$) that there exists an exponent $\nu$ and constants $A^{\prime\prime}_{v}>0$ such that
(3.5) $\langle\|\pi\|^{2}\rangle_{n}^{v}\sim A_{v}^{\prime\prime}n^{2\nu},\qquad v\in V.$
As above, this should hold for $d=4$ with $\nu=\frac{1}{2}$ and subject to the inclusion of the correction term $(\log n)^{1/4}$.
The above three exponents are believed to be related to one another through the so-called Fisher relation
(3.6) $\gamma=\nu(2-\eta).$
It is convenient to work with definitions of critical exponents that do not depend on an assumption of dimensionality, and thus we proceed as follows. Let $G$ be an infinite, connected, quasi-transitive graph with connective constant $\mu$ and fundamental domain $W$. Let $X$ be the set of edges incident to vertices in $W$, and let $\Sigma$ be the set of SAWs on $G$ starting at mid-edges in $X$. We define the function
$Y(x,y)=\sum_{\pi\in\Sigma}\frac{x^{|\pi|}}{|\pi|^{y}},\qquad x>0,\ y\in{\mathbb{R}}.$
(The denominator is interpreted as 1 when $|\pi|=0$.) For fixed $x$, $Y(x,y)$ is non-increasing in $y$. Let $\gamma=\gamma(G)\in[-\infty,\infty]$ be such that
$Y(\mu^{-1},y)\begin{cases}=\infty&\text{if }y<\gamma,\\ <\infty&\text{if }y>\gamma.\end{cases}$
We shall assume that $\infty<\gamma<\infty$. It will be convenient at times to assume more about the number $\sigma_{n}$ of $n$-step SAWs from mid-edges in $X$, namely that there exist constants $C_{i}=C_{i}(W)\in(0,\infty)$ and a slowly varying function $L$ such that
(3.7) $C_{1}L(n)n^{\gamma-1}\mu^{n}\leq\sigma_{n}\leq C_{2}L(n)n^{\gamma-1}\mu^{n},\qquad n\geq 1.$
Let
(3.8) $V(z)=\sum_{n=1}^{\infty}\frac{1}{n^{2z+1}}\langle\|\pi\|^{2}\rangle_{n},\qquad z\in[-\infty,\infty],$
where $\langle\cdot\rangle_{n}$ denotes the uniform average over the set $\Sigma_{n}$ of $n$-step SAWs in $\Sigma$. Thus, $V(z)$ is non-increasing in $z$, and we let $\nu=\nu(G)\in[-\infty,\infty]$ be such that
$V(z)\begin{cases}=\infty&\text{if }z<\nu,\\ <\infty&\text{if }z>\nu.\end{cases}$
Let $\alpha W$ denote the image of $W$ under an automorphism $\alpha\in{\mathcal{A}}$, with incident edges $\alpha X$, and let
$Z_{\alpha}(x)=\sum_{\pi\in\Sigma(\alpha)}x^{|\pi|},$
where $\Sigma(\alpha)$ is the subset of $\Sigma$ containing SAWs ending at mid-edges in $\alpha X$. We assume there exists $\eta=\eta(G)\in[-\infty,\infty]$ such that, for any sequence of automorphisms $\alpha$ satisfying $d_{G}(W,\alpha W)\to\infty$,
(3.9) $Z_{\alpha}(\mu^{-1})d_{G}(W,\alpha W)^{w}\begin{cases}\to 0&\text{if }w<\eta,\\ \to\infty&\text{if }w>\eta.\end{cases}$
The $\eta$ of (3.4) should agree with that defined here when $d=2$.
It is easily seen that the values of $\gamma$, $\eta$, $\nu$ do not depend on the choice of fundamental domain $W$.
We consider now the effect on critical exponents of the local transformation. Let $W_{0}$ be a minimal fundamental domain of $G_{0}:=G$, with incident edge-set $X_{0}:=X$ as above. Write $W_{1}=F(W_{0})$, the set of vertices of the triangles formed by the Fisher transformation at vertices in $W_{0}$, and $X_{1}$ for the set of edges of $G_{1}$ incident to vertices in $W_{1}$. It may be seen that $W_{1}$ is a fundamental domain of $G_{1}$.
Theorem 3.4.
Let $G_{0}$ be an infinite, quasi-transitive, connected, cubic graph. Assume that $|\gamma(G_{0})|<\infty$ and that $\eta(G_{1})$ exists.
  • (a)
    The exponents $\gamma$, $\eta$ of $G_{0}$ and $G_{1}$ are equal.
  • (b)
    Let $\sigma_{n,k}$ be the number of $n$-step SAWs on $G_{k}$ from mid-edges in $X_{k}$. Assume the $\sigma_{n,k}$ satisfy (3.7) for constants $C_{i,k}$ and a common slowly varying function $L$. Then the exponents $\nu$ of $G_{0}$ and $G_{1}$ are equal.
Our final result concerns the effect of the local transformation when applied to just one of the vertex-sets of a bipartite graph. Let $G=(V,E)$ be bipartite with vertex-sets $V_{1}$, $V_{2}$ coloured white and black, respectively. We think of $G$ as a graph together with a colouring $\chi$, and the coloured-automorphism group ${\mathcal{A}}_{\text{\rm c}}={\mathcal{A}}_{\text{\rm c}}(G)$ of the pair $(G,\chi)$ is the set of maps $\phi:V\to V$ which preserve both graph structure and colouring. The coloured graph is quasi-transitive if there exists a finite subset $W\subseteq V$ such that: for all $v\in V$, there exists $\alpha\in{\mathcal{A}}_{\text{\rm c}}$ such that $\alpha v\in W$ and $\chi(v)=\chi(\alpha v)$. As before, such a set $W$ is called a fundamental domain.
Theorem 3.5.
  1. (1)
    Let $G$ be an infinite, connected, bipartite graph with vertex-sets coloured black and white, and suppose that the coloured graph $G$ is quasi-transitive, and every black vertex has degree $3$. Let $\widetilde{G}$ be obtained by applying the local transformation at each black vertex.
    • (a)
      The connective constants $\mu$ and $\widetilde{\mu}$ of $G$ and $\widetilde{G}$, respectively, satisfy $\mu^{-2}=h(\widetilde{\mu}^{-1})$ where $h(x)=xg(x)$.
    • (b)
      Under the corresponding assumptions of Theorem 3.4, the exponents $\gamma$, $\eta$, $\nu$ are the same for $G$ as for $\widetilde{G}$.
  2. (2)
    Let $G$ be an infinite, connected, bipartite graph with vertex-sets coloured black and white, and suppose that the coloured graph $G$ is quasi-transitive, and every vertex has degree $3$. Let $\widetilde{G}$ be obtained by applying the local transformation $\phi_{1}$ at each black vertex; and local transformation $\phi_{2}$ at each white vertex
    • (a)
      The connective constants $\mu$ and $\widetilde{\mu}$ of $G$ and $\widetilde{G}$, respectively, satisfy $\mu^{-2}=h(\widetilde{\mu}^{-1})$ where $h(x)=g_{\phi_{1}}(x)g_{\phi_{2}}(x)$.
    • (b)
      Under the corresponding assumptions of Theorem 3.4, the exponents $\gamma$, $\eta$, $\nu$ are the same for $G$ as for $\widetilde{G}$.
Theorem 3.5(a) implies an exact value of a connective constant that does not appear to have been noted previously. Take $G={\mathbb{H}}$, the hexagonal lattice with connective constant $\mu=\sqrt{2+\sqrt{2}}\approx 1.84776$, see [3].
The proofs of Theorems 3.3-3.4 and 3.5 are found in Sections 4 and 5, respectively.
4. Proof of Theorems 3.3-3.4
Proof of Theorem 3.3.
Let $G=(V,E)$ be an infinite, connected, quasi-transitive, cubic graph. The graph $G_{1}=\phi(G)$ is also quasi-transitive. It suffices for part (a) to show that the connective constants $\mu_{k}$ of the $G_{k}$ satisfy
(4.1) $g(\mu_{1}^{-1})=\mu_{0}^{-1}.$
Let $W$ be a minimal fundamental domain of $G$, and let $X$ be the subset of $E$ comprising all edges incident to vertices in $W$. Write $W_{1}=\phi(W)$, the set of vertices of the triangles formed by the Fisher transformation at vertices in $W$, and $X_{1}$ for the set of edges of $G_{1}$ incident to vertices in $W_{1}$. It may be seen that $W_{1}$ is a fundamental domain of $G_{1}$.
It is convenient to work with SAWs that start and end at mid-edges. Note that the mid-edges of $E$ (respectively, $X$) may be viewed as mid-edges of $E_{1}$ (respectively, $X_{1}$).
Let $G_{0}:=G$. For $k=0,1$, the partition functions of SAWs on $G_{k}$ are the polynomials
$Z_{k}(x)=\sum_{\pi\in\Sigma_{k}}x^{|\pi|},\qquad x>0,$
where the sum is over the set $\Sigma_{k}$ of SAWs starting at mid-edges of $X_{k}$. Similarly, we set
$Z_{1}^{*}(x)=\sum_{\pi\in\Sigma_{1}^{*}}x^{|\pi|},$
where the sum is over the set $\Sigma_{1}^{*}$ of SAWs on $G_{1}$ starting at mid-edges of $X$ and ending at mid-edges of $E$. For $k=0,1$,
(4.2) $Z_{k}(x)\begin{cases}<\infty&\text{if }x<\mu_{k}^{-1},\\ =\infty&\text{if }x>\mu_{k}^{-1}.\end{cases}$
The following basic argument formalizes a method known already in the special case of the hexagonal lattice, see for example [7, 10]. Since $\Sigma_{1}^{*}\subseteq\Sigma_{1}$, we have
(4.3) $Z_{1}^{*}(x)\leq Z_{1}(x).$
Let $N(\pi)$ be the number of endpoints of a SAW $\pi\in\Sigma_{1}$ that are mid-edges of $E$. The set $\Sigma_{1}$ may be partitioned into three sets.
  • (a)
    If $N(\pi)=2$, then $\pi$ contributes to $Z_{1}^{*}$.
  • (b)
    $\pi$ may be a walk within a single gadget.
  • (c)
    If (b) does not hold and $N(\pi)\leq 1$, any endpoint not in $E$ may be moved by finitely many steps (indeed, at most $|\phi(v)|$, where $|\phi(v)|$ is the number of vertices in the gadget $\phi(v)$) along $\pi$ to obtain a shorter SAW in $\Sigma_{1}^{*}$.
By considering the numbers of SAWs in each subcase of (c), we find that
(4.4) $Z_{1}(x)\leq\left[1+\left(\sum_{i=1}^{|\phi(v)|}(Dx)^{i}\right)\right]^{2}Z_{1}^{*}(x)+|E(\phi(v))||W|\left(\sum_{i=1}^{|\phi(v)|}(Dx)^{i}\right).$
where the last term corresponds to case (b) and $|E(\phi(v))|$ is the number of edges in the gadget $\phi(v)$, and $D$ is the maximal vertex degree in $\phi(v)$. By (4.3)–(4.4),
$Z_{1}(x)<\infty\quad\Leftrightarrow\quad Z_{1}^{*}(x)<\infty,$
so that, by (4.2),
(4.5) $Z_{1}^{*}(x)\begin{cases}<\infty&\text{if }x<\mu_{1}^{-1},\\ =\infty&\text{if }x>\mu_{1}^{-1}.\end{cases}$
With a SAW in $\Sigma_{1}^{*}$ we associate a SAW in $\Sigma_{0}$ by shrinking each gadget $\phi(v)$ to a vertex.
Let $r$ be the number of SAW’s in a gadget $\phi(v)$ joining two outer vertices; by Definition 3.1(c), $r$ is independent of the two specific outer vertices chosen. Each $n$-step SAW in $\Sigma_{0}$ arises thus from $r^{n}$ SAWs in $\Sigma_{1}^{*}$. Therefore,
(4.6) $Z_{0}(g(x))=Z_{1}^{*}(x),$
and (4.1) follows by (4.2) and (4.5). ∎
Proof of Theorem 3.4.
Throughout, let $G_{0}:=G$ and $G_{1}:=\phi(G)$, and let $W_{k},X_{k},\Sigma_{k}$ be as in Section 3. Let $\Sigma_{1}^{\ast}$ be the set of SAWs on $G_{1}$ that start at mid-edges of $X_{0}$ and end at mid-edges of $E_{0}$ (viewed as mid-edges of $E_{1}$). For $n\geq 0$, write $\Sigma_{n,k}$ (resp. $\Sigma_{n,1}^{\ast}$) for the subset of $\Sigma_{k}$ (resp. $\Sigma_{1}^{\ast}$) containing $n$-step SAWs, and let $\sigma_{n,k}:=|\Sigma_{n,k}|$ and $\sigma_{n,1}^{\ast}:=|\Sigma_{n,1}^{\ast}|$.
Let $N:=|V(\phi(v))|$ be the number of vertices of a gadget (independent of $v$ by Definition 3.1(2)). Let $\Delta:=\mathrm{diam}(\phi(v))$ and $D:=\max_{u\in V(\phi(v))}\deg_{\phi(v)}(u)$ (both independent of $v$). Note that any SAW segment inside a gadget connecting two ports has length between $2$ and $N$ (in the mid-edge convention of Section 2).
(a) Equality of $\gamma$ and $\eta$. For $k=0,1$ define
$Y_{k}(x,y)\;=\;\sum_{\pi\in\Sigma_{k}}\frac{x^{|\pi|}}{|\pi|^{\,y}},\qquad Y_{1}^{\ast}(x,y)\;=\;\sum_{\pi\in\Sigma_{1}^{\ast}}\frac{x^{|\pi|}}{|\pi|^{\,y}},\qquad x>0,\ y\in\mathbb{R},$
with the convention that the denominator is $1$ when $|\pi|=0$.
Step 1: $Y_{1}^{\ast}(x,y)<\infty\iff Y_{1}(x,y)<\infty$ for fixed $(x,y)$. Since $\Sigma_{1}^{\ast}\subseteq\Sigma_{1}$, we have $Y_{1}^{\ast}(x,y)\leq Y_{1}(x,y)$.
Conversely, consider $\pi\in\Sigma_{1}\setminus\Sigma_{1}^{\ast}$. Either: (i) $\pi$ is entirely contained in a single gadget (there are only finitely many such walks, of bounded length $\leq N$), or (ii) $\pi$ visited at least one mid-edge in $E_{0}$ and at least one endpoint of $\pi$ is a mid-edge not in $E_{0}$. In case (ii), move each endpoint that is not in $E_{0}$ along $\pi$ to the first mid-edge of $E_{0}$ encountered when traversing $\pi$ from that endpoint. This removes at most $N$ visited vertices per endpoint, hence produces a (strictly) shorter SAW $\pi^{\dagger}\in\Sigma_{1}^{\ast}$ with $|\pi|-2N\leq|\pi^{\dagger}|\leq|\pi|$.
Moreover, for a fixed $\pi^{\dagger}$, the number of possible $\pi$ that reduce to it by this procedure is bounded by a constant depending only on $(N,D)$ (each endpoint extension has at most $\sum_{j=0}^{N}(Dx)^{j}$ choices at weight $x$, and there are two endpoints). Since $|\pi|$ and $|\pi^{\dagger}|$ differ by at most $2N$, we also have for fixed $y\in\mathbb{R}$ a constant $C_{y}<\infty$ such that $|\pi|^{-y}\leq C_{y}\,|\pi^{\dagger}|^{-y}$. It follows that for each fixed $(x,y)$,
$Y_{1}(x,y)\;\leq\;C(x,y)\,Y_{1}^{\ast}(x,y)\;+\;C^{\prime}(x,y),$
where $C^{\prime}(x,y)<\infty$ accounts for the finitely many SAWs contained in a single gadget. Therefore,
(4.7) $Y_{1}^{\ast}(x,y)<\infty\iff Y_{1}(x,y)<\infty.$
Step 2: compare $Y_{1}^{\ast}(x,y)$ and $Y_{0}(g(x),y)$. Fix an $n$-step SAW $\omega\in\Sigma_{n,0}$. Contracting each gadget of $G_{1}$ to a vertex maps every $\pi\in\Sigma_{1}^{\ast}$ to a unique $\omega\in\Sigma_{0}$; conversely, expanding each visit to a vertex of $\omega$ by a SAW segment inside the corresponding gadget yields SAWs in $\Sigma_{1}^{\ast}$. Let $a_{m}:=\sigma^{\phi(v)}_{m}(e_{1},e_{2})$ be the number of $m$-step gadget SAWs (in the mid-edge convention) connecting two ports, so that
$g(x)=\sum_{m\geq 0}a_{m}x^{m},\qquad a_{0}=a_{1}=0,\qquad a_{m}=0\text{ for }m>N.$
Let $b_{n,m}$ be the coefficient of $x^{m}$ in $g(x)^{n}$, i.e. $g(x)^{n}=\sum_{m}b_{n,m}x^{m}$. Then $b_{n,m}\neq 0$ only when $2n\leq m\leq Nn$, and the total contribution to $Y_{1}^{\ast}(x,y)$ from all $\pi\in\Sigma_{1}^{\ast}$ that contract to this fixed $\omega$ equals
$T_{n}(x,y)\;:=\;\sum_{m=2n}^{Nn}b_{n,m}\,\frac{x^{m}}{m^{\,y}}.$
Since $2n\leq m\leq Nn$, we have
$(2n)^{-y}\wedge(Nn)^{-y}\;\leq\;m^{-y}\;\leq\;(2n)^{-y}\vee(Nn)^{-y}.$
Hence, with
$c_{y}:=2^{-y}\wedge N^{-y},\qquad c_{y}^{\prime}:=2^{-y}\vee N^{-y},$
we obtain the bounds
$c_{y}\,\frac{g(x)^{n}}{n^{\,y}}\;\leq\;T_{n}(x,y)\;\leq\;c_{y}^{\prime}\,\frac{g(x)^{n}}{n^{\,y}}.$
Summing $T_{n}(x,y)$ over all $\omega\in\Sigma_{n,0}$ and then over $n\geq 1$ gives
(4.8) $c_{y}\,\widetilde{Y}_{0}(g(x),y)\;\leq\;\widetilde{Y}_{1}^{\ast}(x,y)\;\leq\;c_{y}^{\prime}\,\widetilde{Y}_{0}(g(x),y),$
where $\widetilde{Y}$ denotes the series with the (harmless) $n=0$ term removed. In particular,
(4.9) $Y_{1}^{\ast}(x,y)<\infty\iff Y_{0}(g(x),y)<\infty.$
By Theorem 3.3, $\mu_{0}^{-1}=g(\mu_{1}^{-1})$. Combining (4.7) and (4.9) at $x=\mu_{1}^{-1}$ yields
$Y_{1}(\mu_{1}^{-1},y)<\infty\iff Y_{0}(\mu_{0}^{-1},y)<\infty,$
and therefore $\gamma(G_{1})=\gamma(G_{0})$.
Step 3: equality of $\eta$. For $\alpha\in A(G_{0})$, and $k\in\{0,1\}$ define
$Z_{\alpha,k}(x)=\sum_{\pi\in\Sigma_{k}(\alpha)}x^{|\pi|}$
where $\Sigma_{k}(\alpha)$ is the subset of $\Sigma_{k}$ containing SAWs ending at mid-edges in $\alpha X_{k}$. Define $Z_{\alpha,1}^{\ast}(x)$ analogously but with SAWs on $G_{1}$ from mid-edges of $X_{0}$ to mid-edges of $\alpha X_{0}$ (both viewed in $G_{1}$). As in the proof of Theorem 3.3, endpoint-adjustment inside gadgets gives a two-sided bound
(4.10) $Z_{\alpha,1}^{\ast}(x)\;\leq\;Z_{\alpha,1}(x)\;\leq\;C(x)\,Z_{\alpha,1}^{\ast}(x),$
for a finite constant $C(x)$ depending only on the gadget and $x$. Moreover, shrinking gadgets yields the exact identity
(4.11) $Z_{\alpha,1}^{\ast}(x)\;=\;Z_{\alpha,0}(g(x)).$
Evaluating at $x=\mu_{1}^{-1}$ and using $\mu_{0}^{-1}=g(\mu_{1}^{-1})$ gives $Z_{\alpha,1}^{\ast}(\mu_{1}^{-1})=Z_{\alpha,0}(\mu_{0}^{-1})$.
Finally, graph distances in $G_{0}$ and $G_{1}$ are comparable up to multiplicative constants: there exists $\kappa\in(0,\infty)$ (depending only on the gadget) such that for all $\alpha$,
$d_{G_{0}}(W_{0},\alpha W_{0})\;\leq\;d_{G_{1}}(W_{1},\alpha W_{1})\;\leq\;\kappa\,d_{G_{0}}(W_{0},\alpha W_{0}).$
Using (4.10)–(4.11) and the above distance comparison in the definition (3.9) of $\eta$, we obtain $\eta(G_{1})=\eta(G_{0})$ (under the assumed existence of $\eta(G_{1})$). This completes the proof of part (a).
(b) Equality of $\nu$ under (3.7). For $k=0,1$, let
$V_{k}(z)\;=\;\sum_{n=1}^{\infty}\frac{1}{n^{2z+1}}\,\Big\langle\|\pi\|_{k}^{\,2}\Big\rangle_{n,k},$
as in (3.8), and define
(4.12) $V_{1}^{\ast}(z)\;:=\;\sum_{n=1}^{\infty}\frac{1}{n^{2z+1}}\,\frac{1}{\sigma_{n,1}}\sum_{\pi\in\Sigma_{n,1}^{\ast}}\|\pi\|_{1}^{\,2}.$
(Equivalently, $V_{1}^{\ast}(z)=\sum_{n\geq 1}n^{-(2z+1)}(\sigma_{n,1}^{\ast}/\sigma_{n,1})\langle\|\pi\|_{1}^{2}\rangle_{n,1}^{\ast}$.)
Step 4: $V_{1}(z)<\infty\iff V_{1}^{\ast}(z)<\infty$ for each finite $z$. Clearly $V_{1}^{\ast}(z)\leq V_{1}(z)$. For the reverse implication,
Recall that
$V_{1}(z)=\sum_{n\geq 1}\frac{1}{n^{2z+1}}\Big\langle\|\pi\|_{1}^{2}\Big\rangle_{n,1}=\sum_{n\geq 1}\frac{1}{n^{2z+1}\sigma_{n,1}}\sum_{\pi\in\Sigma_{n,1}}\|\pi\|_{1}^{2},$
and, by (4.12),
$V_{1}^{\ast}(z)=\sum_{n\geq 1}\frac{1}{n^{2z+1}}\frac{\sigma_{n,1}^{\ast}}{\sigma_{n,1}}\Big\langle\|\pi\|_{1}^{2}\Big\rangle_{n,1}^{\ast}=\sum_{n\geq 1}\frac{1}{n^{2z+1}\sigma_{n,1}}\sum_{\pi\in\Sigma_{n,1}^{\ast}}\|\pi\|_{1}^{2}.$
Hence
(4.13) $V_{1}(z)=V_{1}^{\ast}(z)+R(z),\qquad R(z):=\sum_{n\geq 1}\frac{1}{n^{2z+1}\sigma_{n,1}}\sum_{\pi\in\Sigma_{n,1}\setminus\Sigma_{n,1}^{\ast}}\|\pi\|_{1}^{2}.$
(i) Walks contained in a single gadget give a finite additive constant. Let $\mathcal{G}$ be the set of SAWs in $\Sigma_{1}$ that are entirely contained in a single gadget. Since the walk starts at a mid-edge in $X_{1}$, only finitely many gadgets intersect the fundamental domain $W_{1}$, and each gadget is finite. Therefore $\mathcal{G}$ is finite, and its total contribution
$C_{z}^{\prime}:=\sum_{\pi\in\mathcal{G}}\frac{\|\pi\|_{1}^{2}}{|\pi|^{2z+1}\sigma_{|\pi|,1}}$
is finite for each fixed finite $z$.
(ii) Trimming map to $\Sigma_{1}^{\ast}$ with bounded multiplicity. For $\pi\in\Sigma_{n,1}\setminus\Sigma_{n,1}^{\ast}$ that is not contained in a single gadget, at least one endpoint of $\pi$ is not a mid-edge of $E_{0}$. Move each endpoint along $\pi$ to the first mid-edge of $E_{0}$ encountered, deleting the initial segment at that endpoint. This produces a shorter SAW $\tau(\pi)\in\Sigma_{m,1}^{\ast}$ with
$m=|\tau(\pi)|\in\{n-j:\ 0\leq j\leq 2N\}.$
Moreover,
$K:=\Bigl(1+\sum_{i=1}^{N}D^{i}\Bigr)^{2}<\infty$
such that every $\pi^{\ast}\in\Sigma_{m,1}^{\ast}$ has at most $K$ preimages (each endpoint can be extended by at most $N$ steps, each step has at most $D$ choices; there are two endpoints) under $\tau$.
Also, trimming changes each endpoint by at most $N$ vertices within a gadget, hence changes the endpoint distance by at most an additive constant. In particular,
$\|\pi\|_{1}\leq\|\tau(\pi)\|_{1}+2M.$
Since $|\tau(\pi)|\geq 1$ implies $\|\tau(\pi)\|_{1}\geq 1$, we have
(4.14) $\|\pi\|_{1}^{2}\leq(1+2M)^{2}\,\|\tau(\pi)\|_{1}^{2}.$
(iii) Comparing the normalizing factors using (3.7). Fix $j\in\{0,\dots,2N\}$ and finite $z$. For $m\geq 1$,
(4.15) $\frac{1}{(m+j)^{2z+1}\sigma_{m+j,1}}=\Bigl(\frac{m}{m+j}\Bigr)^{2z+1}\frac{\sigma_{m,1}}{\sigma_{m+j,1}}\cdot\frac{1}{m^{2z+1}\sigma_{m,1}}.$
By (3.7) (i.e. (3.7)) for $k=1$,
$\sigma_{n,1}\asymp L(n)\,n^{\gamma-1}\mu_{1}^{\,n},$
so for fixed $j$,
$\frac{\sigma_{m,1}}{\sigma_{m+j,1}}\leq\frac{C_{2,1}}{C_{1,1}}\,\frac{L(m)}{L(m+j)}\,\Bigl(\frac{m}{m+j}\Bigr)^{\gamma-1}\,\mu_{1}^{-j}.$
Since $L$ is slowly varying, $L(m)/L(m+j)$ is bounded in $m$ for each fixed $j$, and $\frac{m}{m+j}\in[1/(1+j),1]$ so the power term is bounded as well. Therefore the product
$A_{z,j}:=\sup_{m\geq 1}\Bigl(\frac{m}{m+j}\Bigr)^{2z+1}\frac{\sigma_{m,1}}{\sigma_{m+j,1}}<\infty,$
and (4.15) yields
(4.16) $\frac{1}{(m+j)^{2z+1}\sigma_{m+j,1}}\leq A_{z,j}\cdot\frac{1}{m^{2z+1}\sigma_{m,1}}.$
(iv) Bounding $R(z)$ by $V_{1}^{\ast}(z)$. Using (4.14), bounded multiplicity $K$, and (4.16), we obtain
$R(z)$ $\leq C_{z}^{\prime}+\sum_{j=0}^{2N}\sum_{m\geq 1}\frac{1}{(m+j)^{2z+1}\sigma_{m+j,1}}\sum_{\begin{subarray}{c}\pi\in\Sigma_{m+j,1}\setminus\Sigma_{m+j,1}^{\ast}\\ |\pi|-|\tau(\pi)|=j\end{subarray}}\|\pi\|_{1}^{2}$
$\leq C_{z}^{\prime}+\sum_{j=0}^{2N}\sum_{m\geq 1}\frac{1}{(m+j)^{2z+1}\sigma_{m+j,1}}\sum_{\pi^{\ast}\in\Sigma_{m,1}^{\ast}}K\,(1+2N)^{2}\|\pi^{\ast}\|_{1}^{2}$
$\leq C_{z}^{\prime}+\sum_{j=0}^{2N}K\,(1+2N)^{2}A_{z,j}\sum_{m\geq 1}\frac{1}{m^{2z+1}\sigma_{m,1}}\sum_{\pi^{\ast}\in\Sigma_{m,1}^{\ast}}\|\pi^{\ast}\|_{1}^{2}$
$=C_{z}^{\prime}+C_{z}\,V_{1}^{\ast}(z),$
where $C_{z}:=\sum_{j=0}^{2N}K(1+2N)^{2}A_{z,j}<\infty$. Combining this with (4.13) gives the explicit bound
(4.17) $V_{1}(z)\leq(1+C_{z})\,V_{1}^{\ast}(z)+C_{z}^{\prime}.$
Since also $V_{1}^{\ast}(z)\leq V_{1}(z)$, we conclude that for each fixed finite $z$,
$V_{1}(z)<\infty\iff V_{1}^{\ast}(z)<\infty,$
which is (4.18).
Therefore
(4.18) $V_{1}(z)\;<\infty\iff V_{1}^{\ast}(z)\;<\infty.$
Step 5: $V_{1}^{\ast}(z)<\infty\iff V_{0}(z)<\infty$ for each finite $z$. Fix $n\geq 1$ and $\omega\in\Sigma_{n,0}$. Let $\mathcal{E}(\omega)$ be the family of SAWs $\pi\in\Sigma_{1}^{\ast}$ that contract to $\omega$. As above, for each $m$ the number of $\pi\in\mathcal{E}(\omega)$ with $|\pi|=m$ equals $b_{n,m}$ (the coefficient of $x^{m}$ in $g(x)^{n}$), and $b_{n,m}=0$ unless $2n\leq m\leq Nn$.
Since endpoints of $\pi$ and $\omega$ correspond to the same two original mid-edges, the graph distances satisfy $\|\omega\|_{0}\leq\|\pi\|_{1}\leq\kappa\|\omega\|_{0}$ for some $\kappa<\infty$ depending only on the gadget. Hence, for fixed $\omega$,
$\sum_{\pi\in\mathcal{E}(\omega)}\frac{\|\pi\|_{1}^{\,2}}{|\pi|^{2z+1}\,\sigma_{|\pi|,1}}\asymp\|\omega\|_{0}^{\,2}\sum_{m=2n}^{Nn}b_{n,m}\,\frac{1}{m^{2z+1}\,\sigma_{m,1}},$
where $\asymp$ hides multiplicative constants depending only on $(z,\kappa)$.
Now apply (3.7) (for $k=1$): for all $m\geq 1$,
$\frac{1}{\sigma_{m,1}}\asymp\frac{\mu_{1}^{-m}}{L(m)\,m^{\gamma-1}}.$
Therefore,
$\sum_{m=2n}^{Nn}b_{n,m}\,\frac{1}{m^{2z+1}\,\sigma_{m,1}}\asymp\sum_{m=2n}^{Nn}b_{n,m}\,\frac{\mu_{1}^{-m}}{L(m)\,m^{2z+\gamma}}.$
Since $L$ is slowly varying and $m\in[2n,Nn]$, the values $L(m)$ are comparable to $L(n)$ up to a constant (depending on $N$). Also $m^{-(2z+\gamma)}$ is comparable to $n^{-(2z+\gamma)}$ up to a constant depending on $(z,N)$. Hence
$\sum_{m=2n}^{Nn}b_{n,m}\,\frac{\mu_{1}^{-m}}{L(m)\,m^{2z+\gamma}}\asymp\frac{1}{L(n)\,n^{2z+\gamma}}\sum_{m=2n}^{Nn}b_{n,m}\,\mu_{1}^{-m}.$
But $\sum_{m}b_{n,m}\mu_{1}^{-m}=g(\mu_{1}^{-1})^{n}=\mu_{0}^{-n}$ by Theorem 3.3. Thus,
$\sum_{\pi\in\mathcal{E}(\omega)}\frac{\|\pi\|_{1}^{\,2}}{|\pi|^{2z+1}\,\sigma_{|\pi|,1}}\asymp\|\omega\|_{0}^{\,2}\,\frac{\mu_{0}^{-n}}{L(n)\,n^{2z+\gamma}}.$
Summing over $\omega\in\Sigma_{n,0}$ and then over $n\geq 1$ yields
$V_{1}^{\ast}(z)\asymp\sum_{n\geq 1}\frac{1}{\sigma_{n,0}}\frac{1}{n^{2z+1}}\sum_{\omega\in\Sigma_{n,0}}\|\omega\|_{0}^{\,2}\;=\;V_{0}(z),$
where we used again (3.7) for $k=0$ to identify $\sigma_{n,0}^{-1}n^{-(2z+1)}\asymp\mu_{0}^{-n}(L(n)n^{2z+\gamma})^{-1}$. Consequently, for finite $z$,
(4.19) $V_{1}^{\ast}(z)<\infty\iff V_{0}(z)<\infty.$
Combining (4.18) and (4.19), we conclude that $V_{1}(z)<\infty\iff V_{0}(z)<\infty$ for all finite $z$, hence $\nu(G_{1})=\nu(G_{0})$. ∎
5. Proof of Theorem 3.5
Proof of Theorem 3.5.
We prove part (2). Part (1) follows by taking $\phi_{2}$ to be the trivial transformation (so that $g_{\phi_{2}}(x)=x$ in the mid-edge convention), and by noting that the argument below never uses that white vertices have degree $3$ except when defining $g_{\phi_{2}}$.
Setup. Let $G=(V,E)$ be an infinite, connected, bipartite graph with color classes $V_{\mathrm{black}}$ and $V_{\mathrm{white}}$, and suppose $G$ is quasi-transitive as a colored graph. Assume every vertex has degree $3$. Let $G_{\mathrm{e}}$ be obtained from $G$ by applying the local transformation $\phi_{1}$ at each black vertex and $\phi_{2}$ at each white vertex. Let $W\subset V$ be a minimal fundamental domain for the colored automorphism group, and let $X\subset E$ be the set of edges incident to vertices of $W$. Let $W_{\mathrm{e}}$ be the set of vertices of $G_{\mathrm{e}}$ lying in gadgets over vertices of $W$, and let $X_{\mathrm{e}}$ be the set of edges of $G_{\mathrm{e}}$ incident to vertices in $W_{\mathrm{e}}$. As before, we work with SAWs starting/ending at mid-edges, and the mid-edges of $E$ (resp. $X$) are viewed as mid-edges of $E_{\mathrm{e}}$ (resp. $X_{\mathrm{e}}$).
For $n\geq 0$, let $s_{n}$ be the number of $n$-step SAWs on $G_{\mathrm{e}}$ starting at mid-edges in $X_{\mathrm{e}}$. Let $c_{n}$ be the number of $n$-step SAWs on $G_{\mathrm{e}}$ starting at mid-edges in $X$ and ending at mid-edges in $E$ (both viewed inside $G_{\mathrm{e}}$). Clearly,
(5.1) $c_{n}\leq s_{n}\qquad(n\geq 0).$
Step 1: $s_{n}$ and $c_{n}$ have the same exponential growth rate. Let
$N:=\max\{|V(\phi_{1}(v))|,\ |V(\phi_{2}(v))|\},$
[…]
Every $n$-step SAW counted by $s_{n}$ falls into one of two classes:
  • it is entirely contained in a single gadget (a finite graph);
  • otherwise, it visits at least one mid-edge belonging to $E$.
In the first case $n\leq N$. In the second case, trim each endpoint that is not a mid-edge of $E$ along the walk to the first encountered mid-edge of $E$. This removes at most $N$ visited vertices per endpoint, hence at most $2N$ in total, producing a shorter SAW counted by $c_{n-i}$ for some $i\in\{0,1,\dots,2N\}$.
Conversely, given a walk counted by $c_{n-i}$, the number of ways to extend its endpoints by $i$ additional vertices (inside gadgets) is bounded by $(i+1)D^{i}$. Therefore there exists a constant $C_{0}<\infty$ (depending only on the gadgets and $W$) such that, with the convention $c_{m}=0$ for $m<0$,
(5.2) $s_{n}\leq\sum_{i=0}^{2N}(i+1)D^{i}\,c_{n-i}\qquad(n\geq N+1).$
Combining (5.1) and (5.2) yields
$\limsup_{n\to\infty}s_{n}^{1/n}=\limsup_{n\to\infty}c_{n}^{1/n}.$
Since $G_{\mathrm{e}}$ is quasi-transitive, the connective constant $\mu_{\mathrm{e}}$ exists and equals this common growth rate. In particular, the power series
$C(x):=\sum_{n\geq 0}c_{n}x^{n}$
has radius of convergence $\mu_{\mathrm{e}}^{-1}$.
Step 2: even SAWs on $G$ and the bivariate partition function. Let $\mathcal{E}$ be the set of SAWs on $G$ starting at mid-edges in $X$, and let $\mathcal{E}_{\mathrm{even}}\subset\mathcal{E}$ be those of even length. For $\pi\in\mathcal{E}$, write $|\pi_{\mathrm{black}}|$ (resp. $|\pi_{\mathrm{white}}|$) for the number of black (resp. white) vertices visited by $\pi$; then $|\pi|=|\pi_{\mathrm{black}}|+|\pi_{\mathrm{white}}|$.
For $x,y>0$, define
$Z(x,y):=\sum_{\pi\in\mathcal{E}}x^{|\pi_{\mathrm{black}}|}y^{|\pi_{\mathrm{white}}|},\qquad Z_{\mathrm{even}}(x,y):=\sum_{\pi\in\mathcal{E}_{\mathrm{even}}}x^{|\pi_{\mathrm{black}}|}y^{|\pi_{\mathrm{white}}|}.$
Clearly $Z_{\mathrm{even}}(x,y)\leq Z(x,y)$. Let $\Delta:=\sup_{v\in V}\deg(v)<\infty$ (here $\Delta=3$). Every odd-length SAW in $\mathcal{E}$ is obtained from an even-length SAW by appending one extra step at one endpoint; the number of choices is bounded by $\Delta$, and the one-step SAWs contribute a finite constant depending on $|X|$. Consequently, for all $x,y>0$,
(5.3) $Z(x,y)-Z_{\mathrm{even}}(x,y)\ \leq\ (\Delta x+\Delta y)\,\bigl(|X|+Z_{\mathrm{even}}(x,y)\bigr),$
and hence
(5.4) $Z(x,y)<\infty\iff Z_{\mathrm{even}}(x,y)<\infty.$
Since $G$ is bipartite, any even-length SAW alternates colors and satisfies $|\pi_{\mathrm{black}}|=|\pi_{\mathrm{white}}|=|\pi|/2$, and therefore
$Z_{\mathrm{even}}(x,y)=\sum_{\pi\in\mathcal{E}_{\mathrm{even}}}(xy)^{|\pi|/2}=\sum_{m\geq 0}\sigma_{2m}\,(xy)^{m},$
where $\sigma_{n}$ is the number of $n$-step SAWs in $\mathcal{E}$. Thus the radius of convergence of $Z_{\mathrm{even}}(x,y)$ as a function of $t:=xy$ equals $\mu^{-2}$, where $\mu=\mu(G)$ is the connective constant of $G$. In particular,
(5.5) $Z(x,y)<\infty\iff xy<\mu^{-2}.$
Step 3: substitution identity and the connective-constant equation. Let
$g_{1}(x):=g_{\phi_{1}}(x),\qquad g_{2}(x):=g_{\phi_{2}}(x),$
where each $g_{i}$ is defined as in (3.1) (Definition 3.2) for the corresponding gadget. Consider SAWs counted by $c_{n}$, i.e. SAWs in $G_{\mathrm{e}}$ that start at a mid-edge of $X$ and end at a mid-edge of $E$. Contracting each gadget of $G_{\mathrm{e}}$ to its original vertex yields a map from such SAWs to SAWs $\pi\in\mathcal{E}$ on $G$. Conversely, given $\pi\in\mathcal{E}$, each time $\pi$ visits a black vertex it uses two incident edges, and the portion of the lifted walk inside the corresponding black gadget is any SAW connecting the two corresponding incident mid-edges; the total contribution of all choices is exactly $g_{1}(x)$. Similarly each visited white vertex contributes a factor $g_{2}(x)$. Since $\pi$ is self-avoiding, these gadget-choices occur in disjoint gadgets and hence multiply. Therefore we obtain the exact identity
(5.6) $C(x)=\sum_{n\geq 0}c_{n}x^{n}\;=\;\sum_{\pi\in\mathcal{E}}g_{1}(x)^{|\pi_{\mathrm{black}}|}\,g_{2}(x)^{|\pi_{\mathrm{white}}|}\;=\;Z\bigl(g_{1}(x),g_{2}(x)\bigr).$
By (5.5), we have
$C(x)<\infty\iff g_{1}(x)\,g_{2}(x)<\mu^{-2}.$
Hence the radius of convergence of $C(x)$ is the unique positive solution $\rho$ of
$g_{1}(\rho)\,g_{2}(\rho)=\mu^{-2}.$
Since the radius of $C(x)$ is $\mu_{\mathrm{e}}^{-1}$ by Step 1, this yields
$\mu^{-2}=g_{\phi_{1}}(\mu_{\mathrm{e}}^{-1})\,g_{\phi_{2}}(\mu_{\mathrm{e}}^{-1}),$
which is Theorem 3.5(2)(a).
For part (1)(a), take $\phi_{2}$ to be the trivial transformation; in the mid-edge convention, traversing an untransformed vertex contributes one visited vertex, hence $g_{\phi_{2}}(x)=x$, and the identity above becomes $\mu^{-2}=x\,g_{\phi_{1}}(x)$ with $x=\mu_{\mathrm{e}}^{-1}$.
Step 4: invariance of critical exponents (parts (1)(b) and (2)(b)). Under the corresponding assumptions of Theorem 3.4, one repeats the arguments of Section 4 with the contraction/expansion map used in (5.6) (and the basic even/odd comparison (5.4)) to compare the finiteness of the analogues of $Y(\mu^{-1},y)$, $Z_{\alpha}(\mu^{-1})$, and $V(z)$ for $G$ and $G_{\mathrm{e}}$. The gadget lengths are uniformly bounded above and below, so the polynomial weights in $|\pi|$ are affected only by multiplicative constants as in the proof of Theorem 3.4. This implies that the exponents $\gamma,\eta$ are preserved, and that $\nu$ is preserved under the additional hypothesis (3.7) with a common slowly varying function. ∎
6. Example
Example 6.1 (Complete-graph gadget $K_{N}$).
Fix an integer $N\geq 3$. Define a local transformation $\phi^{(N)}$ as follows: each degree-$3$ vertex $v$ is replaced by a copy of the complete graph $K_{N}$ with three distinguished (outer) vertices $w_{1},w_{2},w_{3}$, and the three edges incident to $v$ are attached to $w_{1},w_{2},w_{3}$ (one per edge). Since $\mathrm{Aut}(K_{N})\cong S_{N}$ contains permutations acting transitively on $\{w_{1},w_{2},w_{3}\}$ while preserving this set, $\phi^{(N)}$ satisfies Definition 3.1.
Let $g_{N}(x)=g_{\phi^{(N)}}(x)$ be as in (3.1). Consider SAWs in $K_{N}$ that start at the mid-edge of the edge attached to $w_{1}$ and end at the mid-edge attached to $w_{2}$. In the mid-edge convention, such a walk visits a sequence of vertices
$w_{1}=v_{1},v_{2},\dots,v_{m}=w_{2},$
with all $v_{i}$ distinct. Since $K_{N}$ is complete, every such sequence is a SAW. Hence, for each $m\in\{2,3,\dots,N\}$ the number of $m$-step gadget SAWs equals the number of ordered choices of the (m2)(m-2) internal vertices from the (N2)(N-2) available vertices $V(K_{N})\setminus\{w_{1},w_{2}\}$, namely
$a_{m}=(N-2)(N-3)\cdots(N-m+1)=\frac{(N-2)!}{(N-m)!}.$
Therefore
(6.1) $g_{N}(x)=\sum_{m=2}^{N}a_{m}x^{m}=\sum_{m=2}^{N}\frac{(N-2)!}{(N-m)!}\,x^{m}=x^{2}\sum_{k=0}^{N-2}(N-2)_{k}\,x^{k},$
where $(N-2)_{k}=(N-2)(N-3)\cdots(N-1-k)$ is the falling factorial.
By Theorem 3.3, if $G$ is cubic with connective constant $\mu$ and $G_{1}=\phi^{(N)}(G)$ has connective constant $\mu_{1}$, then $\mu^{-1}=g_{N}(\mu_{1}^{-1})$. In particular, $\mu_{1}^{-1}$ is the unique solution $x\in(0,1)$ of $g_{N}(x)=\mu^{-1}$.
See Figure 6.1 for an example of a $K_{N}$ gadget with $N=7$.
[…]
Figure 6.1. A local transformation consisting of a $K_{7}$ gadget. In this case, $g_{7}(x)=x^{2}+5x^{3}+20x^{4}+60x^{5}+120x^{6}+120x^{7}$.
Example 6.2 (Generalized Fisher transformations).
Let $H$ be a finite connected graph with two distinct distinguished vertices $v_{1},v_{2}$. Let $H^{(1)},H^{(2)},H^{(3)}$ be three disjoint copies of $H$, and write $v_{1}^{(i)},v_{2}^{(i)}$ for the copies of $v_{1},v_{2}$ in $H^{(i)}$. Form a graph $\widetilde{H}$ by identifying
$v_{2}^{(i)}\equiv v_{1}^{(i+1)}\qquad(i\in\mathbb{Z}/3\mathbb{Z}).$
Let the three ports (outer vertices) be $w_{i}:=v_{1}^{(i)}$ for $i=1,2,3$. Then $\widetilde{H}$ is a gadget: the cyclic permutation of the three copies induces automorphisms acting transitively on $\{w_{1},w_{2},w_{3}\}$. Replacing every degree-$3$ vertex of a cubic graph $G$ by $\widetilde{H}$ (attaching the three incident edges to $w_{1},w_{2},w_{3}$) defines a local transformation, denoted $\phi^{H}$. When $H$ is a single edge, $\phi^{H}$ is the usual Fisher transformation.
To describe the corresponding series $g_{H}(x):=g_{\phi^{H}}(x)$, let $H^{\prime}$ be obtained from $H$ by attaching one extra leaf-edge at each of $v_{1}$ and $v_{2}$, and let $e_{1},e_{2}$ be these two new edges. Define the two-terminal SAW series
$f_{H}(x):=\sum_{\pi\in\Sigma_{H^{\prime},e_{1},e_{2}}}x^{|\pi|}.$
Then
(6.2) $g_{H}(x)=f_{H}(x)+x^{-1}\,f_{H}(x)^{2}.$
Indeed, a gadget-SAW from the port attached to $w_{1}$ to that attached to $w_{2}$ either avoids $w_{3}$ (hence lies in a single copy of $H$ and contributes $f_{H}$), or visits $w_{3}$ (hence splits into two such walks, with $w_{3}$ counted twice, giving the factor $x^{-1}$).
Consequently, if $G$ has connective constant $\mu$ and $\phi^{H}(G)$ has connective constant $\mu_{H}$, then by Theorem 3.3,
$\mu^{-1}=g_{H}(\mu_{H}^{-1}),\qquad\text{and }\ \mu_{H}^{-1}\in(0,1)\ \text{is the unique solution of }\ g_{H}(x)=\mu^{-1}.$
Subexamples.
  1. (1)
    If $H$ is the path on $N\geq 2$ vertices with endpoints $v_{1},v_{2}$, then $f_{H}(x)=x^{N}$ and
    $g_{H}(x)=x^{N}+x^{2N-1}.$
  2. (2)
    If $H$ is the cycle $C_{N}$ ($N\geq 3$) and the graph-distance between $v_{1},v_{2}$ along the shorter arc is $k$, then
    $f_{H}(x)=x^{k+1}+x^{N-k+1},\qquad g_{H}(x)=f_{H}(x)+x^{-1}f_{H}(x)^{2}.$
[…]
Figure 6.2. A generalized Fisher transformation: (a) a cycle graph $H$ on $N=7$ vertices with two distinguished vertices $v_{1},v_{2}$ spaced $k=2$ edges apart and (b) the gadget $\widetilde{H}$ constructed according to Example 6.2.
Example 6.3 (Compositions of local transformations).
Let $G$ be a cubic graph. Suppose $\phi_{1}$ is a local transformation such that $G_{1}:=\phi_{1}(G)$ is again cubic, and let $\phi_{2}$ be a local transformation on $G_{1}$ with $G_{2}:=\phi_{2}(G_{1})$. Write $g_{i}(x):=g_{\phi_{i}}(x)$ and $\mu_{i}:=\mu(G_{i})$. By Theorem 3.3,
$\mu_{0}^{-1}=g_{1}(\mu_{1}^{-1}),\qquad\mu_{1}^{-1}=g_{2}(\mu_{2}^{-1}),$
hence
$\mu_{0}^{-1}=(g_{1}\circ g_{2})(\mu_{2}^{-1}).$
Equivalently, $\mu_{2}^{-1}\in(0,1)$ is the unique solution of $(g_{1}\circ g_{2})(x)=\mu_{0}^{-1}$.
More generally, for a sequence $\phi_{1},\dots,\phi_{r}$ such that $G_{i}:=\phi_{i}(G_{i-1})$ is cubic for $1\leq i\leq r$, one has
$\mu_{0}^{-1}=(g_{1}\circ\cdots\circ g_{r})(\mu_{r}^{-1}),\qquad g_{\phi_{r}\circ\cdots\circ\phi_{1}}=g_{1}\circ\cdots\circ g_{r}.$
Remark 6.4 (Iterating a fixed transformation).
Assume $\phi$ is a local transformation such that $G_{n}:=\phi^{n}(G)$ is cubic for all $n\geq 0$. Let $x_{n}:=\mu(G_{n})^{-1}$ and $g(x):=g_{\phi}(x)$. Then Theorem 3.3 gives
$x_{n-1}=g(x_{n})\qquad(n\geq 1),$
so $x_{n}=g^{-1}(x_{n-1})$.
Since $g$ is a polynomial with nonnegative coefficients and no constant/linear term (in particular $g(0)=g^{\prime}(0)=0$), it is strictly increasing on $(0,1]$ and hence invertible on its image.
Assume in addition that $g$ is strictly convex on $(0,1)$ and that $g(1)\geq 1$. Then h(x):=g(x)xh(x):=g(x)-x is strictly convex with $h(0)=0$, $h^{\prime}(0)=-1$, and $h(1)\geq 0$, so there exists a unique $x_{\ast}\in(0,1]$ such that $g(x_{\ast})=x_{\ast}$. Moreover, $(x_{n})$ is monotone (increasing if $x_{0}<x_{\ast}$ and decreasing if $x_{0}>x_{\ast}$) and converges to $x_{\ast}$.

### v2

@@
 Abstract.
-We study self-avoiding walks (SAWs) on infinite quasi-transitive cubic graphs under local transformations that replace each degree-$3$ vertex by a finite, symmetric three-port gadget.
-To each gadget we associate a two-port SAW generating function $g(x)$, defined by counting SAWs that enter and exit the gadget through prescribed ports.
-Our first main result shows that, if $G$ is cubic and $G_{1}=\phi(G)$ is obtained by applying the local transformation at every vertex, then the connective constants $\mu(G)$ and $\mu(G_{1})$ satisfy the functional relation
-$\mu(G)^{-1}=g\bigl(\mu(G_{1})^{-1}\bigr).$
-We next consider critical exponents defined via susceptibility-type series that do not rely on an ambient Euclidean dimension, and prove that the exponents $\gamma$ and $\eta$ are invariant under local transformations; moreover $\nu$ is invariant under a standard regularity hypothesis on SAW counts (a common slowly varying function).
-Our second set of results concerns bipartite graphs, where the local transformation is applied to one colour class (or to both classes, possibly with different gadgets).
-In this setting we obtain an analogous relation
-$\mu(G)^{-2}=h\bigl(\mu(G_{\mathrm{e}})^{-1}\bigr),$
-with $h(x)=xg(x)$ when only one class is transformed and $h(x)=g_{\phi_{1}}(x)\,g_{\phi_{2}}(x)$ when both are transformed.
-We further present explicit families of examples, including replacing each degree-3 vertex by a complete-graph gadget $K_{N}$.
+Despite its elementary definition, the self-avoiding walk (SAW) poses notoriously hard enumerative problems: exact connective constants are known for only a handful of infinite graphs, notably the honeycomb lattice [3].
+We establish a general substitution principle for SAWs on infinite connected quasi-transitive cubic graphs under port-transitive vertex replacements, where each degree-$3$ vertex is replaced by a fixed finite three-port gadget.
+Writing $g(x)$ for the associated two-port SAW series, we prove that for $G_{1}=\phi(G)$,
+$\mu(G)^{-1}=g\bigl(\mu(G_{1})^{-1}\bigr),$
+equivalently $\mu(G_{1})^{-1}$ is the unique solution $x\in(0,1)$ of $g(x)=\mu(G)^{-1}$, thereby extending the Fisher-triangle relation of Grimmett–Li to arbitrary symmetric three-port gadgets.
+We also obtain the corresponding identity for bipartite graphs when one or both colour classes are transformed, and show that the critical exponents $\gamma$ and $\eta$ (and $\nu$ under a standard regularity hypothesis) are invariant.
+For explicit gadget families, including complete-graph gadgets $K_{N}$ and Fisher-type constructions, these identities turn base graphs with known $\mu$ into infinite families of new quasi-transitive graphs whose connective constants are determined exactly as the unique roots of explicit algebraic equations.
 Key words and phrases:
@@
 1. Introduction
-A self-avoiding walk (SAW) on a graph $G$ is a path that visits no vertex more than once.
-SAWs were introduced in the chemical theory of polymerization (see Flory [5]) and have since been central objects in probability, combinatorics, and statistical physics; see, for example, Madras and Slade [13].
-For an infinite quasi-transitive graph $G$, the number $\sigma_{n}$ of $n$-step SAWs grows exponentially, and the associated exponential growth rate
-$\mu=\mu(G)=\lim_{n\to\infty}\sigma_{n}^{1/n}$
-is the connective constant of $G$. Exact values of connective constants are known only in a small number of nontrivial cases.
-This paper studies how SAWs and their asymptotic parameters behave under local transformations of cubic graphs.
-Informally, a local transformation replaces every degree-$3$ vertex by a fixed finite gadget with three distinguished outer vertices (ports) attached to the three incident edges, subject to a symmetry condition ensuring that the three ports are indistinguishable from the point of view of the gadget (Definition 3.1).
-A canonical example is the Fisher transformation, which replaces each vertex by a triangle; Fisher-type transformations play an important role in exact mappings between two-dimensional lattice models (e.g.
-Ising/dimer/vertex models) and have also been used in the SAW literature.
-Connective constants under local transformations.
-Given a local transformation $\phi$, we associate to the gadget a two-port generating function $g(x)$ (Definition 3.2), defined by counting SAWs that enter the gadget through one port-edge and exit through another, while otherwise staying inside the gadget.
-Our first main theorem (Theorem 3.3) shows that this local data determines the connective constant of the transformed graph: if $G$ is infinite, connected, quasi-transitive and cubic, and $G_{1}=\phi(G)$ is obtained by applying $\phi$ at every vertex, then
-(1.1) $\mu(G)^{-1}=g\bigl(\mu(G_{1})^{-1}\bigr).$
-Equivalently, $\mu(G_{1})^{-1}$ is the unique positive solution of $g(x)=\mu(G)^{-1}$.
-This functional relation may be viewed as a substitution identity for SAW partition functions, and it extends the familiar relation arising from the Fisher transformation.
-Critical exponents.
-Beyond exponential growth, SAWs are expected to exhibit universal power-law corrections, captured by critical exponents.
-Rather than assuming an ambient Euclidean dimension, we work with graph-theoretic definitions of the exponents $\gamma,\eta,\nu$ via the finiteness of natural SAW generating series (Section 3).
-Our second main theorem (Theorem 3.4) proves that local transformations preserve these exponents under mild hypotheses: $\gamma$ and $\eta$ are invariant, and $\nu$ is invariant assuming the SAW counts satisfy a standard slowly varying correction (3.7) with a common slowly varying function.
-Transforming one colour class of a bipartite graph.
-Our third main result concerns bipartite graphs.
-Let $G$ be an infinite, connected, quasi-transitive bipartite graph with colour classes (black/white), and suppose at least one class is cubic.
-Let $G_{\mathrm{e}}$ be obtained by applying a local transformation to all vertices in one class (or, in the fully cubic case, applying possibly different local transformations to the two classes).
-Theorem 3.5 shows that the connective constants satisfy
-(1.2) $\mu(G)^{-2}=h\bigl(\mu(G_{\mathrm{e}})^{-1}\bigr),$
-with $h(x)=xg(x)$ if only one class is transformed, and $h(x)=g_{\phi_{1}}(x)\,g_{\phi_{2}}(x)$ if both classes are transformed.
-This yields new exact connective constants when combined with known exact values.
-The special case in which each cubic vertex is replaced by a triangle is the Fisher transformation.
-The effect of the Fisher transformation on self-avoiding walks was analyzed by Grimmett and Li [6], who derived an explicit functional relation between the connective constants of $G$ and its Fisher transform and proved invariance of the principal critical exponents under iteration.
-The present paper extends [6] from Fisher triangles to general local transformations in which each degree-$3$ vertex is replaced by an arbitrary finite three-port gadget satisfying a natural symmetry condition (Definition 3.1).
-Under this general framework we obtain corresponding functional relations for connective constants (Theorems 3.3 and 3.5) and prove invariance of the critical exponents under verifiable hypotheses (Theorem 3.4), including the case where the transformation is applied to only one colour class of a bipartite graph (Theorem 3.5).
-.
-Further examples.
-We include explicit families of gadgets beyond Fisher triangles.
-In particular, replacing each cubic vertex by a complete graph $K_{N}$ with three designated outer vertices yields an explicit polynomial $g_{N}(x)$ (Example 6.1), and hence an explicit implicit relation between the connective constants of $G$ and $\phi^{(N)}(G)$ for any $N\geq 3$.
-When $N=3$, this corresponds to Fisher transformation.
-See Figure 1.1 for local replacements of a degree-3 vertex by a complete graph $K_{3}$ and $K_{4}$.
-[…]
-Figure $1.$ Local replacement at a degree-$3$ vertex: (a) the original star, (b) Fisher transformation (triangle gadget), and (c) replacement by a $K_{4}$ gadget.
-See also Definition 3.1 and Figure 3.1 in the draft.
-A useful class of examples comes from vertex-transitive TLF-planar cubic graphs.
-Following Renault ([14]), planarity plus vertex-transitivity imposes strong geometric symmetry on embeddings: the action of graph automorphisms can (locally) be represented by isometries of the ambient geometry (spherical, Euclidean, or hyperbolic).
-Let $G$ be a transitive, TLF-planar cubic graph, and let $\langle a,b,c\rangle$ be its type-vector, i.e., the sizes of the three faces incident to a vertex, with the convention $1/\infty=0$. Define
-$f(G):=\frac{1}{a}+\frac{1}{b}+\frac{1}{c}.$
-By [14] (see also [8]), $G$ admits a planar TLF embedding by regular polygons in the sphere / Euclidean plane / hyperbolic plane according as $f(G)>\frac{1}{2}$, $f(G)=\frac{1}{2}$, or $f(G)<\frac{1}{2}$,
-In the Euclidean case $f(G)=\frac{1}{2}$, the only possible type-vectors are $\langle 6,6,6\rangle$, $\langle 3,12,12\rangle$, $\langle 4,8,8\rangle$, $\langle 4,6,12\rangle$, and $\langle 4,4,\infty\rangle$; see Figure 1.2.
-$\langle 6,6,6\rangle$ $\langle 4,8,8\rangle$ $\langle 3,12,12\rangle$
-$\langle 4,6,12\rangle$ $\langle 4,4,\infty\rangle$
-Figure 1.2. Local regular-polygon patches for the five Euclidean type-vectors of transitive TLF-planar cubic graphs.
+A self-avoiding walk (SAW) on a graph is a finite path that visits no vertex more than once.
+Introduced as a model for polymer chains, SAWs have become central objects in probability, combinatorics, and statistical physics; see, for example, [5, 12].
+Let $G$ be an infinite connected quasi-transitive graph, and let $\sigma_{n}(v)$ denote the number of $n$-step SAWs starting at $v$.
+A classical theorem of Hammersley implies that the limit
+$\mu(G)=\lim_{n\to\infty}\sigma_{n}(v)^{1/n}$
+exists and is independent of $v$; the constant $\mu(G)$ is the connective constant of $G$.
+Exact values (or explicit characterizations) of $\mu(G)$ are known only in a small number of nontrivial cases.
+Local vertex replacements and a substitution principle.
+We study SAWs on cubic graphs under a class of vertex-replacement operations.
+A local transformation $\phi$ replaces each degree-$3$ vertex by a fixed finite three-port gadget whose three ports attach to the incident edges, subject to a port-transitivity condition on the gadget automorphism group (Definition 3.1).
+The classical example is the Fisher transformation, where each vertex is replaced by a triangle.
+Grimmett and Li [6] analysed SAWs under the Fisher transformation, obtaining an explicit functional relation between connective constants and proving invariance of key critical exponents under iterated Fisher transforms.
+The aim of the present paper is to establish a general exact substitution framework that (i) extends the Fisher-triangle analysis to arbitrary port-transitive three-port gadgets, and (ii) treats the bipartite setting in which the transformation is applied to one colour class, or to both classes with possibly different gadgets.
+The resulting identities are global (they describe radii of convergence and critical exponents) but are driven by finite, explicitly computable local data.
+A two-port gadget series and connective constants. Fix $\phi$, and let $H$ denote its gadget. We associate to $H$ a two-port generating function
+(1.1) $g(x)=\sum_{\pi}x^{|\pi|},$
+where $\pi$ ranges over SAWs that enter $H$ through one prescribed port-edge and exit through another while otherwise remaining inside $H$ (Definition 3.2).
+Our first main theorem (Theorem 3.3) shows that this purely local series governs the connective constant after transforming every vertex: if $G$ is infinite, connected, quasi-transitive, and cubic, and $G_{1}=\phi(G)$, then
+(1.2) $\mu(G)^{-1}=g\bigl(\mu(G_{1})^{-1}\bigr),$
+equivalently $\mu(G_{1})^{-1}$ is the unique $x\in(0,1)$ solving $g(x)=\mu(G)^{-1}$.
+Critical exponents on quasi-transitive graphs.
+To avoid assuming an ambient Euclidean dimension, we adopt graph-theoretic definitions of critical exponents $\gamma,\eta,\nu$ via finiteness of natural susceptibility-type series (Section 3).
+Our second main theorem (Theorem 3.4) proves that local transformations preserve these exponents under mild hypotheses: $\gamma$ and $\eta$ are invariant in general, and $\nu$ is invariant provided the SAW counts satisfy a standard regularity hypothesis with a common slowly varying correction.
+Bipartite transformations.
+If $G$ is bipartite and the transformation is applied to one colour class (or, in the fully cubic case, to both classes with possibly different gadgets), we obtain an analogous relation (Theorem 3.5):
+$\mu(G)^{-2}=h\bigl(\mu(G_{\mathrm{e}})^{-1}\bigr),$
+with $h(x)=xg(x)$ in the one-class case and $h(x)=g_{\phi_{1}}(x)\,g_{\phi_{2}}(x)$ in the two-class case.
+These formulas turn known values of $\mu(G)$ into explicit equations for $\mu(G_{\mathrm{e}})$.
+Examples and computations.
+Section 6 develops explicit gadget families and worked base-graph examples.
+We emphasize complete-graph gadgets $K_{N}$ (yielding closed-form polynomials $g_{N}$), a mixed bipartite transformation on the hexagonal lattice (white $K_{4}$, black Fisher triangles), and an iterated transformation example that exhibits convergence of the sequence of connective constants to a gadget-determined fixed point.
 Organization.
-Section 2 collects basic notation.
-Local transformations and the associated generating function $g(x)$ are introduced in Section 3, where we also state Theorems 3.3–3.5.
-The proofs of Theorems 3.3–3.4 and Theorem 3.5 appear in Sections 4 and 5, respectively.
-We conclude with explicit examples in Section 6.
+Section 2 introduces notation.
+Section 3 defines local transformations and the associated series $g(x)$ and states Theorems 3.33.5.
+Proofs appear in Sections 4 and 5.
+Section 6 contains explicit examples, computations, and remarks on iteration.
 2. Notation
@@
 where $\sigma_{n}(v,w)$ is the number of $n$-step SAWs with endpoints $v$, $w$.
-It is known under certain circumstances that the generating functions $Z_{v,w}$ have radius of convergence $\mu^{-1}$ (see [13, Cor.
+It is known under certain circumstances that the generating functions $Z_{v,w}$ have radius of convergence $\mu^{-1}$ (see [12, Cor.
 3.2.6]), and it is believed that there exists an exponent $\eta$ and constants $A^{\prime}_{v}>0$ such that
@@
 ∎
-6. Example
-Example 6.1 (Complete-graph gadget $K_{N}$).
+6. Examples
+This section collects explicit gadget families and worked base-graph examples illustrating Theorems 3.3 and 3.5.
+Many natural quasi-transitive cubic graphs are planar (e.g.
+the hexagonal lattice), while others are not (e.g.
+the $3$-regular tree); our substitution identities apply uniformly in all cases.
+Example 6.1 (Complete-graph gadgets $K_{N}$).
 Fix an integer $N\geq 3$.
@@
 where $(N-2)_{k}=(N-2)(N-3)\cdots(N-1-k)$ is the falling factorial.
-By Theorem 3.3, if $G$ is cubic with connective constant $\mu$ and $G_{1}=\phi^{(N)}(G)$ has connective constant $\mu_{1}$, then $\mu^{-1}=g_{N}(\mu_{1}^{-1})$.
-In particular, $\mu_{1}^{-1}$ is the unique solution $x\in(0,1)$ of $g_{N}(x)=\mu^{-1}$.
+By Theorem 3.3, if $G$ is cubic with connective constant $\mu$ and $G_{1}=\phi^{(N)}(G)$ has connective constant $\mu_{1}$, then
+$\mu^{-1}=g_{N}(\mu_{1}^{-1}),$
+and $\mu_{1}^{-1}$ is the unique solution $x\in(0,1)$ of $g_{N}(x)=\mu^{-1}$.
 See Figure 6.1 for an example of a $K_{N}$ gadget with $N=7$.
+Remark 6.2 (A worked base graph: the $3$-regular tree).
+Let $T_{3}$ denote the $3$-regular tree. Then $\mu(T_{3})=2$ (indeed $\sigma_{n}=3\cdot 2^{n-1}$). For $G_{1}=\phi^{(N)}(T_{3})$, Theorem 3.3 yields $g_{N}(\mu_{1}^{-1})=\tfrac{1}{2}$. For instance,
+$\mu_{1}\approx 2.16730\ (N=4),\qquad\mu_{1}\approx 2.60068\ (N=5),\qquad\mu_{1}\approx 3.55052\ (N=7),$
+obtained by solving $g_{N}(x)=\tfrac{1}{2}$ for $x\in(0,1)$ and setting $\mu_{1}=1/x$.
+This illustrates how complete-graph gadgets produce explicit, finite polynomial equations for new connective constants.
 […]
 Figure 6.1. A local transformation consisting of a $K_{7}$ gadget. In this case, $g_{7}(x)=x^{2}+5x^{3}+20x^{4}+60x^{5}+120x^{6}+120x^{7}$.
+Example 6.3 (A planar six-vertex “triangle-in-triangle” gadget).
+Let $\phi^{\triangle}$ be the local transformation whose gadget $T$ is the following $6$-vertex graph (see Figure 6.2).
+The gadget has three ports (outer vertices) $w_{1},w_{2},w_{3}$.
+On each side $w_{i}w_{j}$ of the outer triangle we insert one vertex $u_{ij}$, so the boundary edges are
+$w_{1}-u_{12}-w_{2},\qquad w_{2}-u_{23}-w_{3},\qquad w_{3}-u_{31}-w_{1},$
+and we add the three edges forming the inner triangle
+$u_{12}u_{23},\qquad u_{23}u_{31},\qquad u_{31}u_{12}.$
+The dihedral symmetries act transitively on $\{w_{1},w_{2},w_{3}\}$, so $\phi^{\triangle}$ satisfies Definition 3.1.
+Let $g_{\triangle}(x):=g_{\phi^{\triangle}}(x)$ be defined by (3.1).
+A gadget-SAW from the port-edge at $w_{1}$ to that at $w_{2}$ is exactly a simple path from $w_{1}$ to $w_{2}$ inside $T$ (mid-edge convention).
+Enumerating simple paths by their number of visited vertices gives
+$g_{\triangle}(x)=x^{3}+3x^{4}+4x^{5}+2x^{6}.$
+Therefore, for any infinite connected quasi-transitive cubic graph $G$ with connective constant $\mu$ and $G_{1}=\phi^{\triangle}(G)$ with connective constant $\mu_{1}$, Theorem 3.3 yields the exact relation
+$\mu^{-1}=g_{\triangle}\!\bigl(\mu_{1}^{-1}\bigr).$
+Hexagonal lattice. If $G=\mathbb{H}$ is the hexagonal lattice, then $\mu(\mathbb{H})=\sqrt{2+\sqrt{2}}$ [3], so $\mu(\phi^{\triangle}(\mathbb{H}))^{-1}\in(0,1)$ is the unique solution of
+$x^{3}+3x^{4}+4x^{5}+2x^{6}=\bigl(2+\sqrt{2}\bigr)^{-1/2}.$
+(Numerically, this gives $\mu(\phi^{\triangle}(\mathbb{H}))\approx 1.9318771$.)
+AABBCCtriangle-in-triangle gadget
+Figure 6.2. A planar 66-vertex gadget with outer ports $w_{1},w_{2},w_{3}$ and an inner triangle supported on the three sides. For this gadget, $g_{\triangle}(x)=x^{3}+3x^{4}+4x^{5}+2x^{6}$.
+Example 6.4 (A mixed bipartite transformation on the hexagonal lattice).
+Let $H$ be the hexagonal lattice with its natural bipartition $V_{\mathrm{w}}\cup V_{\mathrm{b}}$ (white/black).
+Apply $\phi^{(4)}$ at each white vertex (a $K_{4}$ gadget), and apply the Fisher triangle transformation at each black vertex (i.e.
+a $K_{3}$ gadget).
+Denote the resulting graph by $H_{\mathrm{e}}$.
+For the Fisher triangle one has
+$g_{K_{3}}(x)=x^{2}+x^{3},$
+and for the $K_{4}$ gadget one has $g_{K_{4}}(x)=x^{2}+2x^{3}+2x^{4}$. By Theorem 3.5(2),
+$\mu(H)^{-2}=g_{K_{4}}\bigl(\mu(H_{\mathrm{e}})^{-1}\bigr)\,g_{K_{3}}\bigl(\mu(H_{\mathrm{e}})^{-1}\bigr).$
+Since $\mu(H)^{-2}=(2+\sqrt{2})^{-1}=\tfrac{1}{2}(2-\sqrt{2})$, writing $x:=\mu(H_{\mathrm{e}})^{-1}$ yields the explicit polynomial equation
+(6.2) $x^{4}+3x^{5}+4x^{6}+2x^{7}=\tfrac{1}{2}(2-\sqrt{2}),\qquad x\in(0,1).$
+Thus $\mu(H_{\mathrm{e}})=1/x\approx 1.91529$.
+The two gadgets used here are depicted in Figure 6.3.
+[…]
+Figure 6.3. The two gadgets used in Example 6.4: a Fisher triangle ($K_{3}$) and a $K_{4}$ gadget with ports $A,B,C$.
 Example 6.5 (Generalized Fisher transformations).
@@
     $f_{H}(x)=x^{k+1}+x^{N-k+1},\qquad g_{H}(x)=f_{H}(x)+x^{-1}f_{H}(x)^{2}.$
-[…]
-Figure 6.2.
-A generalized Fisher transformation: (a) a cycle graph $H$ on $N=7$ vertices with two distinguished vertices $v_{1},v_{2}$ spaced $k=2$ edges apart and (b) the gadget $\widetilde{H}$ constructed according to Example 6.2.
 Example 6.6 (Compositions of local transformations).
@@
 $\mu_{0}^{-1}=(g_{1}\circ\cdots\circ g_{r})(\mu_{r}^{-1}),\qquad g_{\phi_{r}\circ\cdots\circ\phi_{1}}=g_{1}\circ\cdots\circ g_{r}.$
+Example 6.7 (A “hub–triangle” gadget and an explicit iteration picture).
+Define a local transformation $\psi$ by replacing each degree-$3$ vertex by the gadget obtained as follows: start from a triangle on outer vertices $w_{1},w_{2},w_{3}$ (the ports), subdivide each of the three triangle edges by inserting a new vertex at its midpoint, and finally add a central vertex connected to the three midpoint vertices.
+(See Figure 6.4.) This gadget is port-transitive by the dihedral symmetry of the construction.
+A direct enumeration of gadget SAWs (in the mid-edge convention of Section 2) yields
+(6.4) $g_{\psi}(x)=x^{3}+4x^{5}+2x^{7}.$
+In particular, $g_{\psi}$ is strictly increasing and strictly convex on $(0,1)$ and satisfies $g_{\psi}(1)=7$.
+If $G_{n}:=\psi^{n}(G)$ remains cubic for all $n\geq 0$ and $x_{n}:=\mu(G_{n})^{-1}$, then Theorem 3.3 gives $x_{n-1}=g_{\psi}(x_{n})$ for $n\geq 1$.
+Since $g_{\psi}$ is strictly convex with $g_{\psi}(0)=0$ and $g_{\psi}^{\prime}(0)=0$, the equation $g_{\psi}(x)=x$ has a unique solution $x_{\ast}\in(0,1)$, and the sequence $(x_{n})$ is monotone and converges to $x_{\ast}$ (cf.
+Remark 6.8 below).
+Numerically, $x_{\ast}\approx 0.6050003$, hence the limiting connective constant is $1/x_{\ast}\approx 1.65289$.
+    AABBCC(a) triangle–hub gadget     $yxx_{\ast}$(b) $y=g_{\psi}(x)$ and $y=x$
+Figure 6.4.
+(a) The triangle–hub gadget used to define the local transformation $\psi$.
+(b) The fixed point $x_{\ast}$ solving $g_{\psi}(x)=x$ for $g_{\psi}(x)=x^{3}+4x^{5}+2x^{7}$.
+So $x_{\ast}=\sqrt{\frac{\sqrt{3}-1}{2}}\approx 0.6050003$, and $\mu(\psi^{n}(G))\longrightarrow\mu_{\ast}:=x_{\ast}^{-1}\approx 1.6528917$.
 Remark 6.8 (Iterating a fixed transformation).
@@
 Moreover, $(x_{n})$ is monotone (increasing if $x_{0}<x_{\ast}$ and decreasing if $x_{0}>x_{\ast}$) and converges to $x_{\ast}$.
+Acknowledgements
+The authors thank the organizers of the conference “Perfectly matched perspectives on statistical mechanics, combinatorics and geometry” (CIRM, 16–20 June 2025) for a stimulating meeting.
+The conference was supported by the National Science Foundation under Grant DMS-2512154.
+B.G.
+was partially supported by a U.S.
+Department of Education GAANN Fellowship.